OpenStudy (anonymous):

Find all the real square roots of 0.0009 Answer a. b. no real solution c. d.

OpenStudy (anonymous):

a.+or-0.3 c.+3 d.+or-0.01

myininaya (myininaya):

\[\sqrt{.0009}=\sqrt{\frac{9}{10000}}=\frac{\sqrt{9}}{\sqrt{10000}}=\frac{\sqrt{3^2}}{\sqrt{10^3}}\] That should be a good hint :)

myininaya (myininaya):

and since it says square roots just do plus or minus in front of that

OpenStudy (anonymous):

I got plus or minus .3

myininaya (myininaya):

oops that was suppose to be 10^4 my bad

myininaya (myininaya):

\[\frac{\sqrt{3^2}}{\sqrt{10^4}}=\frac{3}{10^2}=\frac{3}{100}\]

OpenStudy (apoorvk):

myininaya. i think when they post some thing like this \[x^2 =4\] then we write + or - before sqrt(a) but when it asks \[x = \sqrt{4}\] then we write only +2 as the answer. i donno why, although it means pretty much the same right??

OpenStudy (apoorvk):

i meant sqrt(4) instead of sqrt(a) over there.